arcsine(1/2) = pi/6
because we are looking for angle represented by sinx=1/2.
sin(30) = 1/2 (By looking at Unit Circle or using a calculator), then converting 30 to Radians. The answer is in Radians and is in exact form (There are 6 30s in 180, therefore pi/6 is the first 30).
What is arcsine(1/4) in Radial Exact Form.
I know I get a Right Triangle H=1, Height = 1/4, Length = root(15)/4.
Using a calculator I can get the decimal form ~14 (14.blahblahblah) degrees.
How can I get the exact radial form?

Answer by Alan3354(69443)

arcsine(1/2) = pi/6
because we are looking for angle represented by sinx=1/2.
sin(30) = 1/2 (By looking at Unit Circle or using a calculator), then converting 30 to Radians. The answer is in Radians and is in exact form (There are 6 30s in 180, therefore pi/6 is the first 30).
What is arcsine(1/4) in Radial Exact Form.
I know I get a Right Triangle H=1, Height = 1/4, Length = root(15)/4.
Using a calculator I can get the decimal form ~14 (14.blahblahblah) degrees.
How can I get the exact radial form?
---------------
pi radians = 180 degs
14 degrees --> 14*pi/180 radians